OpenStudy (oneokrock):

please help on how to FOIL after you find the answer to a factoring question.I also don't know what factoring method to use while looking at an equation. Can I just use the long MA table all the time ? example:4x^2+19x-5

OpenStudy (welshfella):

No need to bother with foil expanding factors is easy - use the distributive law (x + 2)((x - 3) = x(x - 3) + 2(x - 3) = x^2 - 3x + 2x - 6 = x^2 - x - 6 just multiply the second parentheses by 3 then by +2

OpenStudy (welshfella):

* multiply the second parentheses by x ( not 3!)

OpenStudy (mww):

The factoring method depends on your situation. For quadratic trinomials (3 terms) there are two ways 4x^2+19x-5 First|middle |last 1) The first involves finding two numbers that multiply to make the PRODUCT of the first and last term (i.e. the coefficient of x^2 and the constant) AND SUM to give the middle term (coefficient of x). P: first x last term: 4 x - 5 = -20 S = 19 So we need two numbers that multiply to make -20 and also add to 19. -1 and 20 are the two numbers we need. Once you have these two numbers you can SPLIT the middle term (19x) into terms with these two as coefficients. So 4x^2+19x-5 becomes 4x^2+20x - x -5 Then group in pairs and factorise 4x^2+20x - x -5 = 4x(x+5) - (x+5) = (4x-1)(x+5) An alternative approach is to use boxes once you have figured out the two special numbers. 

OpenStudy (mww):